OpenStudy (anonymous):

When two atoms of deuterium are fused to form one atom of helium, the total energy evolved is 3.83 * 10^-12 joules. What is the total change in mass in kilograms for this reaction?

OpenStudy (anonymous):

I did this: m = E/c^2 m = 3.83 * 10^-12 J / 9.00 * 10^16 m^2/s^2 m = 4.256 * 10^-28 kg Is this correct?

OpenStudy (aaronq):

it's 10^-29

OpenStudy (anonymous):

how?